OpenStudy (anonymous):

What is the correct first step for solving the equation 5x + 3 = 8? A. Subtract 8 from each side. B. Subtract 5 from each side. C. Subtract 3 from each side. D. Divide each side by 3.

jagr2713 (jagr2713):

C

OpenStudy (anonymous):

Subtract 3 from each side

jagr2713 (jagr2713):

5x+3=8 -3 -3 5x=5 x=5

jagr2713 (jagr2713):

x=1 sorry

OpenStudy (anonymous):

Because if you subtract 3 from itself and 8, it would equal 5x=5 Then you divide 5x too 5 to get x=1

OpenStudy (anonymous):

Y U beat me jagr?

jagr2713 (jagr2713):

lol

OpenStudy (anonymous):

hey ima guest its guggle 113 i changed accounts this is my new one

OpenStudy (anonymous):

Oh, really?

OpenStudy (anonymous):

Why did you change?

OpenStudy (anonymous):

i didnt like my old username

OpenStudy (anonymous):

Makes sense. So, I can see you like Halo.

OpenStudy (anonymous):

yea did u notice the red vs blue reference

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Its been a long time since I seen the series. I stopped watching at season 10

OpenStudy (anonymous):

Hey are you in FLVS?

OpenStudy (anonymous):

FLVS?

OpenStudy (anonymous):

Guess not. FLVS: Florida Virtual School.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i live in pa

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

haha we should probably message eachother rather than be in this guys question

OpenStudy (anonymous):

Yeah, well...

OpenStudy (anonymous):

lol i meant mine

OpenStudy (anonymous):

nvm its my question were good

OpenStudy (anonymous):

Well if you need anymore help with these types of questions, let me know.

OpenStudy (anonymous):

wait do u play xbox

OpenStudy (anonymous):

Used to. I have an Xbox 360 but my younger brother plays on it now. I occasionally get on it to play Halo 3 and Dark Souls. But besides that, I play on PS4 and PS3.

OpenStudy (anonymous):

I am planning on getting an Xbox One soon though, not determined if I will get it though.